OpenStudy (anonymous):

Amy paid $84.17 for a pair of shoes during a 25% off sale. What was the regular price?

OpenStudy (anonymous):

Let the regular price be x. x - 25x/100 = 84.17 Solve for x.

OpenStudy (anonymous):

84.17 = 75% Divide by 75 to give 1% Multiply by 100 to give 100%

OpenStudy (anonymous):

\[\LARGE x(1-25\%)=84.17\]

OpenStudy (anonymous):

So the answer is?

OpenStudy (anonymous):

\[\LARGE x-\frac{25x}{100}=84.17 \] as pratu said... solve for X and you'll have the answer .